jeff is 10 years old and his brother is 2 years old. in how many years will jeff be twice as old as his brother?
please help me for this all problems i have send.thank you so much

Answer by ankor@dixie-net.com(22740)

let j = Jeff's present age
let b = brother's
:
Write an equation for the given statement
:
jeff is 10 years old and his brother is 2 years old.
in how many years will jeff be twice as old as his brother?
let y = no. of years for this to be true
j + y = 2(b+y)
we know j=10; b=2
10 + y = 2(2+y)
10 + y = 4 + 2y
10 - 4 = 2y - y
6 = y
In 6 yrs J will be twice as old as B
:
Check that
10 + 6 = 2(2 + 6)
16 = 2(8)